Erika's uncle is three time as old as she is now. Four years ago, he was four times as old as she was then. How old is Erika now?
Answer by ankor@dixie-net.com(22740)

Erika's uncle is three time as old as she is now.
Four years ago, he was four times as old as she was then.
How old is Erika now?
:
Let e = Erika's age now
Let u = uncles age now
:
Write an equation for each statement:
:
"Erika's uncle is three time as old as she is now."
u = 3e
:
" Four years ago, he was four times as old as she was then."
u - 4 = 4(e-4)
u - 4 = 4e - 16
u = 4e - 16 + 4
u = 4e - 12
:
How old is Erika now?
Replace u with 3e in the above equation
3e = 4e - 12
12 = 4e - 3e
e = 12 yrs is Erika now
:
:
Check solution in the statement:
(u = 3(12) = 36)
"Four years ago, he was four times as old as she was then."
36 - 4 = 4(12 - 4)
32 = 4(8); confirms our solution